A stationary bomb explodes into 3 pieces .one piece of 2 kg mass moves with a velocity of 8 m/s at right angles to the other piece of mass 1kg moving with a velocity of 12m/s .If the mass of the third piece is 0.5kg ,then the velocity of the third piece is?

Grade:12

2 Answers

Tejasvi dalla
7 years ago
As two masses move at right angles.... So net momentum can be calculates via square sum root....[(2*8)^² + (1*12)^²]^½ = 20Acc. To law of conservation of linear momentum....P initial= P final20= 0.5 vv = 40m/s
Khimraj
7 years ago
Just apply conservation of linear momentum as external force is zero. 0 = 2*8i^ + 1*12j^ +1/2v. so V=40m/sec.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer